OpenStudy (anonymous):

The following is a list of the number of siblings that various students have: 1, 1, 3, 0, 4, 1, 2, 1, 2, 0, 3, 2, 1, 5, 4, 2. What is the mean? Answer (A) 1 (B) 2 (C) 4 (D) 5

OpenStudy (anonymous):

Add all the terms first...

OpenStudy (anonymous):

add them up then divide it by the addens(amount of numbers given)

OpenStudy (anonymous):

1 + 1 + 3 + 4+1+2+1+2+3+2+1+5+4+2 = ??

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

32/16=2

OpenStudy (maheshmeghwal9):

\[\huge{\frac{1}{n} \sum_{x=1}^{n}a_x.}\]

OpenStudy (anonymous):

when u add them all up it = 32 then tale the total number of siblings there are which is 16 and divide 32/16

OpenStudy (goformit100):

(D) 5

OpenStudy (maheshmeghwal9):

where n is number of terms looks beautiful summation no need to go deep @LAGIRL69 :-)

OpenStudy (anonymous):

Thanks @laxplayer <3 ....and Everyone else who Help me understand Algebra 1 :*

OpenStudy (anonymous):

@maheshmeghwal9 ha Thanks

OpenStudy (anonymous):

@goformit100 how did u get 5

OpenStudy (maheshmeghwal9):

:)

OpenStudy (anonymous):

yup @LAGIRL69 <3

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

To be sure is it B or D?

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

Thanks always

OpenStudy (anonymous):

the mean is.. add up all of the numbers then divide it by how many numbers u have

OpenStudy (anonymous):

so its B

OpenStudy (anonymous):

Thank you again :)